OpenStudy (anonymous):

Show that \[\int{du \over \sqrt{1 + u^2}} = \ln|u + \sqrt {1 + u^2}| + c\]

OpenStudy (anonymous):

I used a table but then was later told that I needed to do it by hand

OpenStudy (wasiqss):

ok its gonna be done by substitution method

OpenStudy (wasiqss):

or wait can we do that \[d/du= \ln [u=\sqrt{1+u^2}]=\int\limits_{?}^{?}(1)/\sqrt{1+u^2)}\]

OpenStudy (wasiqss):

?

OpenStudy (anonymous):

Use \[ u=\tan(x)\\ du= \sec^2(x) dx\\ \sqrt{ 1 + u^2} =\sqrt{ 1 +\tan^2(x)}= \sec(x)\\ \frac{du}{\sqrt{ 1 + u^2}}=\frac{\sec^2(x) dx} {\sec(x)}= \sec(x) dx \] You can now finish it from here.